OpenStudy (anonymous):

r=lim 1/2*lim(1+1/n)/1 NEXT STEP: r=1/2*1+0/1 ^^ How do you get that?

OpenStudy (anonymous):

wat is this

OpenStudy (anonymous):

pre calculus limits.

OpenStudy (anonymous):

oh one minute I'll look it u

OpenStudy (anonymous):

Just keep the following fact in mind: limn!1 n 1=n = limn!1 e 1 n logn = e limn!1 1 n logn = e 0 = 1 because limx!1 1 x log x = limx!1 log x x L’H= limx!1 1=x 1 = 0 We’ll also use Striling’s approximation: p 2 n n+1=2 e n n! n n+1=2 e n1 for all n 2 N Problem 1 (3.6). (a) P1 n=1 n+1 n Diverges: limn!1 n+1 n = 1 6= 0. In general, if P n an converges, then limn!1 an = 0 – sometimes this is called the Divergence Test for some reason. (f) P1 n=1 n 4 2n Converges: lim sup n 4 2n 1=n = limn!1 n 4=n 2 = limn!1 (n 1=n) 4 2 = 1=2 < 1. Convergence is by the Root Test. (g) P1 n=1 p n (n)n = P1 n=1 (1)n nn2 Converges: limn!1 1 nn2 = 0, so convergence is by the Alternating Series Test. (h) P1 n=1 n! nn Converges: Observe that, n! nn = 1 2 3 (n 1) n nn = 1 n 2 n 3 n n 1 n n n = = 1 n 2 n 3 n n 1 n n n < 2 n2 Then, convergence is by the Comparison Test – comparing to P1 n=1 2 n2 , which converges by the p-Test. (i) P1 n=1 sin n 2 Diverges: limn!1 sin n 2 doesn’t exist – see (a). (l) P1 n=1 n+2 n3+5 Converges: Could use the Root Test, but I’ll use the Comparison Test. n + 2 n3 + 5 = 1 n3 1 n2 + 2 n3 1 + 5 n3 1 n3 3 6 = 1 2n3 and P1 n=1 1 2n3 converges by the p-Test. 1

OpenStudy (anonymous):

wat happened i copy and paste and thats wat happened

OpenStudy (anonymous):

@Hope_nicole write your equation clearly and legibly.

OpenStudy (anonymous):

\[\lim \frac{ 1 }{ 2 }* \lim \frac{ 1+1/n }{ 1 }\] the next step goes to this:\[\frac{ 1 }{ 2 }*\frac{ 1+0 }{ 1 }\]

OpenStudy (anonymous):

i dont see how you get from the first step to the last step.

OpenStudy (anonymous):

what the the limit to???? limits look like this: \[\lim_{x\to a}f(x)\]

OpenStudy (anonymous):

idk how to write it with the equation thing on here but it is n-->infinity.

OpenStudy (anonymous):

ok. you see. that information is crucial.. otherwise, it is like saying, "I bought __les with __dol__"

OpenStudy (anonymous):

so, is this like this? \[\lim_{n\to\infty}{1\over2}\lim_{n\to\infty}1+{1\over n}\]

OpenStudy (anonymous):

no the last one is\[\frac{ 1+1/n }{ 1}\]

OpenStudy (anonymous):

ok.. remember that anything divided by "1" gives you the same thing

OpenStudy (anonymous):

\[1\times x=x\implies x={x\over 1}\]

OpenStudy (anonymous):

you limit \[ L=\left(\lim_{n\to\infty}{1\over2}\right)\left(\lim_{n\to\infty}1+{1\over n}\right)\\ \] the first paranthesis is independent of "n" so, it will not change. and in the second one, we subtitute the limit \[L={1\over2}\left(1+{1\over\infty}\right)\\ \text{but}\quad{1\over\infty}=0\\ L={1\over2}(1+0)={1\over2}\]